# Align & Invert: Solving Inverse Problems with Diffusion and Flow-based Models via Representational Alignment

### v1

@@ A.1 Proofs
$\widetilde{h}_{t}^{(\mathrm{repa})}=\widetilde{h}_{t}+2\lambda\,J_{\widetilde{G}_{2}}(z_{t})\,J_{\widetilde{G}_{2}}(z_{t})^{\!\top}\,\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{h}_{t}\bigr)+O(\lambda^{2}).$
Let
$A_{t}:=J_{\widetilde{G}_{2}}(z_{t})J_{\widetilde{G}_{2}}(z_{t})^{\top}\;\in\;\mathbb{R}^{ND_{2}\times ND_{2}}.$
Subtracting $\widetilde{h}^{*}$ on both sides yields
$\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}$ $=(\widetilde{h}_{t}-\widetilde{h}^{*})+2\lambda\,A_{t}\,\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{h}_{t}\bigr)+O(\lambda^{2}).$
We now add and subtract $\widetilde{f}(x)$ and $\widetilde{\Phi}\widetilde{h}^{*}$ inside the parentheses:
[…]
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{f}(x)\bigr)$
[…]
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{\Phi}\widetilde{h}^{*}-\widetilde{\Phi}\widetilde{h}_{t}\bigr)+O(\lambda^{2}).$
Grouping terms gives
[…]
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{f}(x)\bigr)$
$\quad+2\lambda A_{t}\widetilde{\Phi}^{\top}\bigl(\widetilde{f}(x)-\widetilde{\Phi}\widetilde{h}^{*}\bigr)+O(\lambda^{2}).$
Taking Euclidean norms and using the triangle inequality yields
$\|\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}\|_{2}$ $\leq\|I-2\lambda A_{t}\widetilde{\Phi}^{\top}\widetilde{\Phi}\|_{2}\,\|\widetilde{h}_{t}-\widetilde{h}^{*}\|_{2}$
$\quad+2\lambda\|A_{t}\widetilde{\Phi}^{\top}\|_{2}\Bigl(\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}+\|\widetilde{f}(x)-\widetilde{\Phi}\widetilde{h}^{*}\|_{2}\Bigr)+O(\lambda^{2}).$
By the definition of $\mathrm{ApproxErr}$,
$\mathrm{ApproxErr}(x,\bar{x})=\frac{1}{N}\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(x)-f_{\mathrm{DINOv2}}^{[n]}(\bar{x})\bigr\|_{2}^{2},$
and in vectorized form this implies
$\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}^{2}=\sum_{n=1}^{N}\bigl\|f_{\mathrm{DINOv2}}^{[n]}(\bar{x})-f_{\mathrm{DINOv2}}^{[n]}(x)\bigr\|_{2}^{2}=N\,\mathrm{ApproxErr}(x,\bar{x}),$
so
$\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}=\sqrt{N}\,\sqrt{\mathrm{ApproxErr}(x,\bar{x})}.$
Similarly, from the definition of $\mathrm{MisREPA}$ we obtain
$\|\widetilde{f}(x)-\widetilde{\Phi}\widetilde{h}^{*}\|_{2}=\sqrt{N}\,\sqrt{\mathrm{MisREPA}(x)}.$
Substituting into the previous inequality gives
$\|\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}\|_{2}\leq\|I-2\lambda A_{t}\Phi^{\top}\Phi\|_{2}\,\|\widetilde{h}_{t}-\widetilde{h}^{*}\|_{2}+2\lambda\sqrt{N}\,\|A_{t}\Phi^{\top}\|_{2}\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right)+O(\lambda^{2}).$
Since $A_{t}=J_{\widetilde{G}_{2}}(z_{t})J_{\widetilde{G}_{2}}(z_{t})^{\top}$ and $\widetilde{\Phi}^{\top}\widetilde{\Phi}$ are positive semidefinite,
$\|A_{t}\|_{2}=\sigma_{\max}(J_{\widetilde{G}_{2}}(z_{t}))^{2},\qquad\|\widetilde{\Phi}^{\top}\widetilde{\Phi}\|_{2}=\sigma_{\max}(\widetilde{\Phi})^{2},$
where $\sigma_{\max}(\cdot)$ denotes the largest singular value. Since $\widetilde{\Phi}=I_{N}\otimes\Phi$, its singular values are exactly those of $\Phi$ (each repeated $N$ times), hence $\|\widetilde{\Phi}^{\top}\widetilde{\Phi}\|_{2}=\sigma_{\max}(\Phi)^{2}$.
Thus, choosing
$0<\lambda<\frac{1}{\sigma_{\max}(J_{\widetilde{G}_{2}}(z_{t}))^{2}\,\sigma_{\max}(\Phi)^{2}}$
ensures
$\|I-2\lambda A_{t}\Phi^{\top}\Phi\|_{2}\leq C_{1}\qquad\text{for some }0<C_{1}<1.$
Letting
$C_{2}:=2\lambda\,\sqrt{N}\,\|A_{t}\Phi^{\top}\|_{2},$
and absorbing the $O(\lambda^{2})$ term into the constants for sufficiently small $\lambda$, we obtain the vectorized bound
$\|\widetilde{h}_{t}^{(\mathrm{repa})}-\widetilde{h}^{*}\|_{2}\;\leq\;C_{1}\,\|\widetilde{h}_{t}-\widetilde{h}^{*}\|_{2}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right).$
Finally, since for any $H\in\mathbb{R}^{k_{1}\times k_{2}}$ we have
$\|H\|_{F}=\|\mathrm{vec}(H)\|_{2},$
the same inequality holds in matrix form with Frobenius norms:
$\|h_{t}^{(\mathrm{REPA)}}-h^{*}\|_{F}\;\leq\;C_{1}\,\|h_{t}-h^{*}\|_{F}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right),$
which is exactly the claim of the proposition.
@@ A.2 Details About Latent DPS + REPA Implementation
We implement Latent DPS + REPA following the procedure in Algorithm 2. A key design choice is the learning rate schedule $\eta_{t}$, for which we consider two adaptive parameterizations:
$\eta(t)=\left\{\begin{aligned} &\frac{\kappa}{\|y-A\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\|_{2}}&&\text{(Adaptive inverse norm)}\\[12.0pt] &\frac{\kappa}{\max\!\left(\tfrac{t}{1-t},\,1\right)}&&\text{(Adaptive SNR-based learning rate)}\end{aligned}\right$.
where $\kappa$ is a tunable scaling factor. Table 3 compares these schedules for $4\times$ super-resolution on ImageNet, showing that the SNR-based strategy consistently achieves better perceptual performance (lower LPIPS) and slightly higher PSNR. For each schedule, $\kappa$ and $\lambda$ are tuned on a small validation set. We use the SNR-based learning rate for all main experiments unless otherwise.
Table 3: Comparison of learning rate strategies for $4\times$ super-resolution on ImageNet.
(a) Latent DPS
LR Type $\kappa$ PSNR ↑ LPIPS ↓
Inverse norm 7.75 26.09 0.274
SNR-based 2.25 26.82 0.238
(b) Latent DPS + REPA
LR Type $\lambda$ $\kappa$ PSNR ↑ LPIPS ↓
Inv norm 0.005 5 25.90 0.235
SNR-based 0.01 2 26.82 0.217
@@
1:flow model $u_{\theta}$ = $G_{1}\circ G_{2}$, measurement $y$, pretrained encoder $f$
@@
7:  $z_{t-1}\leftarrow z_{t}+\Delta t\cdot v$
@@
9:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos((f_{\text{DINOv2}}^{[n]}(y),g_{\phi}(G_{2}(z_{t},t)^{[n]}))$
10:end for
11:return $x_{0}$
@@ A.3 Details About Resample + REPA Implementation
In Table 3, we present an adaptation of the algorithm proposed by song2024solving to the flow-based setting, augmented with the REPA regularizer as described in the methodology section. We note that the original ReSample algorithm employs a three-stage procedure for enforcing data consistency: (i) gradient steps in latent space, similar to Latent DPS; (ii) pixel-space optimization, which is computationally efficient and captures high-level semantics but often leads to blurrier reconstructions; and (iii) latent space optimization as outlined to line 7 of Algorithm 3. In contrast, our adaptation omits the pixel-space stage, as our focus is on maximizing perceptual quality. We found that this modification together with the inclusion of the REPA reguralizer yields sharper and more visually convincing results, while still benefiting from the refinement effect of the final latent-space consistency updates.
Importantly, the resulting algorithm remains consistent with the central principle of ReSample—that initialization strongly impacts the effectiveness of data-consistency updates. In our case, REPA guides the initialization toward reconstructions with stronger perceptual quality, which the subsequent consistency updates then refine further, improving fidelity and sharpness.
Algorithm 3 Latent DPS + REPA Algorithm
1:flow model $u_{\theta}=G_{1}\circ G_{2}$, measurement $y$, pretrained encoder $f$, resample steps $C$, parameter $\gamma$
@@
7:  $z_{t-1}\leftarrow z_{t}+\Delta t\cdot v$
@@
13:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\big((f_{\text{DINO}}^{[n]}(y),\,g_{\phi}(G_{2}(z_{t},t)^{[n]})\big)$
14:end for
@@ A.4 Robustness of DINOv2 Representations to Corruptions
Figure 5 presents how DINOv2 representation similarity changes as the corruption severity increases for the two tasks. Figure 6 further visualizes how the pixel-space appearance of a single image changes across corruption levels. Interestingly, we observe that DINOv2 representations remain significantly more robust to both super-resolution and Gaussian deblurring. Despite substantial visual degradation in pixel space, DINOv2 maintains a strong representation alignment with the original image.
@@
A.5 Enforcing Feature-Level DINOv2 Consistency without Representation Alignment
@@
$z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\left(f_{\text{DINOv2}}^{[n]}(y),\,f_{\text{DINOv2}}^{[n]}\!\big(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\big)\right),$ (17)
While both objectives influence the sampling trajectory, the feature-only variant derives its guidance from the DINOv2 features of the decoded image. In contrast, REPA aligns the diffusion model’s internal representations with DINOv2 features directly at each timestep.
@@
A.6 Additional Qualitative Results
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Figure 8: Qualitative comparison for Box Inpainting. Each row shows: Measurement, pixel-space method DPS, latent method DPS, its REPA-enhanced version, and the Ground Truth.
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Figure 9: Qualitative comparison for $4\times$ super-resolution. Each row shows: Measurement, a baseline method (DPS or DAPS), a latent solver (Latent DPS or ReSample), its REPA-enhanced version, and the Ground Truth.
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth
Measurement DPS Latent DPS Latent DPS + REPA Ground Truth
Measurement Latent DAPS ReSample ReSample + REPA Ground Truth

### v2

@@ A.1 Proofs
$\|z_{t}^{(\mathrm{REPA})}-z^{*}\|_{2}$ $\leq(1-\lambda m_{J}m_{\phi})\,\|z_{t}-z^{*}\|_{2}+C\Bigl(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\Bigr),$ (23)
where the constant $C$ is defined explicitly as
$C:=2\lambda M_{G}M_{\phi}.$
This concludes the proof. ∎
@@ A.2 Details About Latent DPS + Repa Implementation
We implement Latent DPS + Repa following Algorithm 2 and use 1000 sampling steps throughout. We employ an SNR-based learning rate schedule of the form
$\eta(t)=\frac{\kappa}{\max\!\left(\tfrac{t}{1-t},\,1\right)},$
where $\kappa$ is a tunable scaling factor. This schedule accounts for the varying signal-to-noise ratio across diffusion timesteps and provides stable performance in practice. The parameters $\kappa$ and the representation-alignment strength $\lambda$ are tuned on a small validation set with present the best parameters found in Table 2.
Table 2: Hyperparameters used for Latent DPS + Repa. ImageNet uses Gaussian noise $\sigma=0.01$, while FFHQ uses $\sigma=0.05$.
ImageNet FFHQ
Task $\boldsymbol{\kappa}$ $\boldsymbol{\lambda}$ $\boldsymbol{\kappa}$ $\boldsymbol{\lambda}$
Super-resolution 2.0 0.01 0.025 1.25
Gaussian deblurring 0.25 0.05 0.05 0.5
Motion deblurring 0.5 0.01 0.05 0.5
Box inpainting 0.5 0.01 0.025 0.75
@@
1:flow model $u_{\theta}$, measurement $y$, pretrained encoder $f$, proxy representation $c_{\text{proxy}}$, timestep $t_{\text{cutoff}}$
@@
7:  $z_{t-1}\leftarrow z_{t}-\Delta t\cdot v$
@@
9:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos(c_{\text{proxy}}^{[n]},g_{\phi}(\mathrm{DiffEnc}^{[n]}(z_{t},t))$
10:  if $t<t_{\text{cutoff}}$ then
11:   $c_{\text{proxy}}\leftarrow f_{\mathrm{DINOv2}}\!\bigl(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\bigr)$
12:  end if
13:end for
14:return $x_{0}$
@@ A.3 Details About Resample + REPA Implementation
In Table 3, we present an adaptation of the algorithm proposed by Song et al. (2024) to the flow-based setting, augmented with the REPA regularizer as described in the methodology section. We note that the original ReSample algorithm employs a three-stage procedure for enforcing data consistency: (i) gradient steps in latent space, similar to Latent DPS; (ii) pixel-space optimization, which is computationally efficient and captures high-level semantics but often leads to blurrier reconstructions; and (iii) latent space optimization as outlined to line 7 of Algorithm 3. In contrast, our adaptation omits the pixel-space stage, as our focus is on maximizing perceptual quality. We found that this modification together with the inclusion of the REPA reguralizer yields sharper and more visually convincing results, while still benefiting from the refinement effect of the final latent-space consistency updates. The best hyperparameters used for this variant are reported in Table 3. In addition to the parameters $\kappa$ and $\lambda$, which play analogous roles to those in Latent DPS, we also tune the maximum number of inner-loop optimization steps and the corresponding inner learning rate of the Resample procedure (see Algorithm 3).
Table 3: Hyperparameters used for Resampling + Repa. ImageNet uses Gaussian noise $\sigma=0.01$, while FFHQ uses $\sigma=0.05$.
ImageNet FFHQ
Task $\boldsymbol{\kappa}$ $\boldsymbol{\lambda}$ max iters inner lr $\boldsymbol{\kappa}$ $\boldsymbol{\lambda}$ max iters inner lr
Super-resolution 0.05 3.25 150 0.005 0.01 2.25 100 0.0001
Gaussian deblurring 0.075 0.5 300 0.005 0.05 0.75 300 0.00075
Motion deblurring 0.5 0.75 300 0.005 0.05 0.75 100 0.005
Box inpainting 0.025 100 200 0.0005 0.05 0.75 200 0.0005
Algorithm 3 Resample + REPA Algorithm
1:flow model $u_{\theta}$, measurement $y$, pretrained encoder $f$, resample steps $C$, parameter $\gamma$, proxy representation $c_{\text{proxy}}$, timestep $t_{\text{cutoff}}$
@@
7:  $z_{t-1}\leftarrow z_{t}-\Delta t\cdot v$
@@
13:  $z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\big(c_{\text{proxy}}^{[n]},\,g_{\phi}(\mathrm{DiffEnc}^{[n]}(z_{t},t))\big)$
14:  if $t<t_{\text{cutoff}}$ then
15:   $c_{\text{proxy}}\leftarrow f_{\mathrm{DINOv2}}\!\bigl(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\bigr)$
16:  end if
17:end for
@@ A.4 Robustness of DINOv2 Representations to Corruptions
Figure 4 presents how DINOv2 representation similarity changes as the corruption severity increases for the two tasks. Figure 5 further visualizes how the pixel-space appearance of a single image changes across corruption levels. Interestingly, we observe that DINOv2 representations remain significantly more robust to both super-resolution and Gaussian deblurring. Despite substantial visual degradation in pixel space, DINOv2 features maintain a strong alignment with the original image features.
@@
A.5 Selection of Proxy Representation
Since pretrained DINOv2 features are robust to a wide range of common image corruptions, we use the DINOv2 encoding of the measurement as an initial proxy representation, $f_{\mathrm{DINOv2}}(y)$. We find this choice to be effective across all considered inverse problems. In addition, we introduce a cutoff timestep at which the proxy representation is updated to depend on the model’s current denoised estimate, namely $f_{\mathrm{DINOv2}}(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}]))$.
  • Box Inpainting. We construct the proxy representation by combining the observed regions of the measurement with the model’s current denoised estimate:
    $c_{\text{proxy}}=\mathrm{mask}\odot f_{\mathrm{DINOv2}}(y)+(I-\mathrm{mask})\odot f_{\mathrm{DINOv2}}\!\bigl(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\bigr),$ (24)
    where $\mathrm{mask}$ denotes the binary inpainting mask.
  • Super-resolution, Gaussian Deblurring, and Motion Deblurring. For these tasks, we use $f_{\mathrm{DINOv2}}(y)$ as the proxy representation during the early stages of sampling. On the FFHQ dataset, we find it beneficial to switch to the reconstruction-based proxy $f_{\mathrm{DINOv2}}(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}]))$ after $80\%$ of the diffusion steps. On ImageNet, this transition yields only marginal improvements, and we therefore use the measurement-based proxy throughout.
A.6 Experimental Details
We implement the inverse problem operators following the setups of Chung et al. (2023). For deblurring tasks, we follow Zhang et al. (2025) and fix a single realization of the degradation operator to ensure fair and consistent comparisons across methods. All baseline results are obtained using the official implementations provided by Chung et al. (2023) and Zhang et al. (2025).
For DPS, we use the dataset-specific hyperparameters reported in the original paper. For DAPS, we adopt the hyperparameters provided in the official codebase for latent diffusion models. For PSNR and SSIM, we report results obtained by averaging the reconstructed images across five independent runs of the solver and then computing the metrics on the averaged reconstruction.
A.7 Enforcing Feature-Level DINOv2 Consistency without Representation Alignment
@@
$z_{t-1}\leftarrow z_{t-1}+\lambda\nabla_{z_{t}}\sum_{n=1}^{N}\cos\!\left(c^{[n]}_{\text{proxy}},\,f_{\text{DINOv2}}^{[n]}\!\big(\mathcal{D}(\mathbb{E}[z_{0}\mid z_{t}])\big)\right),$ (25)
Note that in this case we can only use $f_{\text{DINOv}2}(y)$ as the proxy term. While both objectives influence the sampling trajectory, the feature-only variant derives its guidance from the DINOv2 features of the decoded image. In contrast, REPA aligns the diffusion model’s internal representations with DINOv2 features directly at each timestep.
@@
A.8 Representation Alignment on an Independently Pretrained Model
The effectiveness of Repa for inverse problems relies on the availability of a mapping $g_{\phi}$ between the internal representations of the diffusion model and the DINOv2 feature space. Such a mapping is naturally available when the diffusion model is pretrained with representation alignment. However, when starting from a diffusion model that has been pretrained independently, this alignment is not present.
To construct the mapping $g_{\phi}$, we finetune the pretrained diffusion model using the representation alignment objective introduced in Eq. 6. However, optimizing the REPA objective in isolation can cause the diffusion model to drift away from the natural image manifold, degrading its denoising and generative performance. To prevent this drift and preserve the learned image prior, we jointly optimize the standard velocity prediction objective $\mathcal{L}_{\mathrm{velocity}}$ (Eq. 3). The resulting finetuning objective is
$\mathcal{L}_{\mathrm{finetune}}:=-\textsc{Repa}(x,h_{t})+\lambda\,\mathcal{L}_{\mathrm{velocity}},$ (26)
where $x$ denotes a clean input image, $h_{t}$ denotes the corresponding internal diffusion representation at noise level $t$, extracted from a fixed intermediate layer of the pretrained diffusion model as defined in 2.2. We set $\lambda=1$ in all experiments.
We evaluate this finetuning strategy on the FFHQ dataset. Specifically, we first train an SiT model without representation alignment for 100k iterations and subsequently perform full finetuning of the model for an additional 20k iterations using the combined objective above to learn the alignment mapping $g_{\phi}$. We then apply Repa to inverse problems using the finetuned model. Table 5 reports the corresponding results, alongside those obtained with a model pretrained with representation alignment.
Across all inverse problems, adding Repa consistently improves perceptual metrics (LPIPS and FID), demonstrating that finetuning with the proposed alignment objective is sufficient to enable effective use of Repa at inference time. While the strongest gains are obtained when the diffusion model is pretrained with representation alignment, the finetuned model exhibits the same qualitative behavior, indicating that alignment can be introduced post hoc without retraining the diffusion model from scratch.
To examine this behavior more closely, we conduct the following analysis. We select a subset of 1000 images from the FFHQ validation set and measure the cosine alignment between the internal representations of the diffusion model and those of the DINOv2 encoder across diffusion timesteps. Figure 7 shows that, although finetuning successfully learns a meaningful alignment mapping, the model pretrained with representation alignment consistently achieves higher alignment values throughout the denoising trajectory. In terms of Definition 4.1, this corresponds to a smaller $\mathrm{MisREPA}$ error, explaining why Repa provides more accurate regularization signals for inverse problems when applied to an alignment-pretrained model.
Table 5: Performance comparison on FFHQ ($\sigma=0.05$) for models trained and finetuned with Repa. Lower is better for LPIPS/FID, higher is better for PSNR/SSIM. Best values per task are shown in bold.
Task Method LPIPS↓ FID↓ PSNR↑ SSIM↑
$4\times$ SR Latent DPS (REPA model) 0.188 56.69 28.99 0.814
Latent DPS + REPA (REPA model) 0.177 51.27 29.12 0.819
Latent DPS (SiT) 0.190 61.36 28.92 0.805
Latent DPS + REPA (SiT finetuned) 0.184 59.22 28.95 0.805
Gaussian Deblur Latent DPS (REPA model) 0.192 60.65 28.15 0.783
Latent DPS + REPA (REPA model) 0.186 55.53 28.21 0.787
Latent DPS (SiT) 0.193 60.72 28.02 0.782
Latent DPS + REPA (SiT finetuned) 0.188 59.76 28.07 0.779
Motion Deblur Latent DPS (REPA model) 0.170 52.14 27.20 0.773
Latent DPS + REPA (REPA model) 0.165 47.18 27.16 0.772
Latent DPS (SiT) 0.171 54.01 27.63 0.779
Latent DPS + REPA (SiT finetuned) 0.167 52.45 27.75 0.780
Figure 7: Cosine alignment between internal diffusion representations and DINOv2 encoder across diffusion timesteps.
A.9 Additional Qualitative Results
Measurement DPS Without REPA With REPA Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption